I need help figuring out how to calculate cost of debt and equity.

Question 2 Key facts and assumptions concerning Organic Grocery at Sept 30, 2018, appear below. Using this information, answer the questions following, as of October 1, 2018. Facts and Assumptions Organic Grocery (WF) Instructions: Yield to maturity on short-term government bonds Yield to maturity on long-term government bonds Coupon rate on WF long term bonds, annual payment Maturity, long term bonds Current bond price Market risk premium Tax rate Estimated company equity beta Stock price per share Number of Equity shares outstanding (millions) Book value of equity (millions) Book value of interest-bearing debt (1 bond; millions S 2.54% 4.54% 7.50% 20 years $1,131.89 6.30% 25.0% 1.00 25.97 681.2 4,965 a. Estimate Organic Grocerys cost of debt. Use the current YTM on debt as an estimate of required return b. Estimate Organic Grocerys cost of equity capital using CAPM c. Estimate Organic Grocerys weighted-average cost of capital. Prepare a table showing the relevant steps & variables clearly labeled. d. SHOW ALL STEPS in arriving at a solution. Clearly label all items used. $ 6,674 Cost of Debt to the Firm: Extra Work Space
Answer #1
a) Pre-tax Cost of the debt to the firm = YTM = 6.32%
Calculation of YTM:
YTM is that discount rate which equates the PV of the cash
flows from the bond with its price.
The cash flows are the maturity value of $1000 receivable
at EOY20 and the annual interest of $75, which is an
annuity.
The YTM has to be found out by trial and error, such that the
PV of cash flows equal the price of the bond.
Using 6%, PV of cash flows = 1000/1.06^20+75*(1.06^20-1)/(0.06*1.06^20) = $    1,172.05
Using 7%, PV of cash flows = 1000/1.07^20+75*(1.07^20-1)/(0.07*1.07^20) = $    1,052.97
The discount rate [YTM}, lies between 6% and 7%.
By simple interpolation, it is
=6+(1172.05-1131.89)/(1172.05-1052.97) = 6.34
Using an online calculator it is 6.32% (This has been used)
After tax cost of debt = 6.32*(1-0.25) = 4.74
b) Cost of equity = Risk free rate+Beta*Market risk premium = 4.54+1*6.30 = 10.84
c) Market Value Weight Component Cost WACC
Value of debt = Book value *113.189% = 6674*113.189% = 7554.23 0.2992 4.74 1.42
Market value of equity = 25.97*681.20 = 17690.76 0.7008 10.84 7.60
Total value 25245.00 9.02
WACC = 9.02%